# Learning to Communicate using Contrastive Learning

### v1

@@ Abstract
Figure 1: Multi-view contrastive learning and CACL, contrastive learning for multi-agent communication. In multi-view learning, augmentations of the original image or “views” are positive samples to contrastively learn features. In CACL, different agents’ views of the same environment states are considered positive samples and messages are contrastively learned as encodings of the state.
@@ 1 Introduction
Communication is a key capability necessary for effective coordination among agents in partially observable environments. In multi-agent reinforcement learning (MARL) (Sutton and Barto, 2018), agents can use their actions to transmit information (Grupen et al., 2020) but continuous or discrete messages on a communication channel (Foerster et al., 2016), i.e., linguistic communication (Lazaridou and Baroni, 2020), are more flexible and powerful because they can convey more complex concepts. To successfully communicate, a speaker and a listener must share a common language with a shared understanding of the symbols being used (Skyrms, 2010; Dafoe et al., 2020). Emergent communication or learning a common protocol (Wagner et al., 2003; Lazaridou and Baroni, 2020), is a thriving research direction but most works focus on simple, single-turn, sender-receiver games (Lazaridou et al., 2018; Chaabouni et al., 2019). In more visually and structurally complex MARL environments (Samvelyan et al., 2019), existing approaches often rely on centralized learning mechanisms by sharing models (Lowe et al., 2017) or gradients (Sukhbaatar et al., 2016).
@@
Inspired by the literature in representation learning that uses different “views” of a data sample (Bachman et al., 2019), for a given trajectory, we propose that an agent’s observation is a “view” of the environment state. Thus, different agents’ messages are encodings of different incomplete “views” of the same underlying state. From this perspective, messages from the same state should be more similar to each other than to those from distant states or other trajectories. We visually show our perspective in Figure 1. We propose Communication Alignment Contrastive Learning (CACL), which each agent uses contrastive learning between sent and received messages to learn to communicate.
@@ 2 Related Work
Learning to coordinate multiple RL agents is a challenging and unsolved task where naively applying single-agent RL algorithms often fails (Foerster et al., 2016). Recent approaches focus on neural network-based agents (Goodfellow et al., 2016) with a message channel to develop a common communication protocol (Lazaridou and Baroni, 2020). To handle issues of non-stationarity, some work focuses on centralized learning approaches that globally share models (Foerster et al., 2016), training procedures (Lowe et al., 2017), or gradients (Sukhbaatar et al., 2016) among agents. This improves coordination and can reduce optimization issues but results are often still sub-optimal in practise (Foerster et al., 2016; Lin et al., 2021) and may violate independence assumptions, effectively modelling the multi-agent scenario as a single agent (Eccles et al., 2019).
This work focuses on independent, decentralized agents and non-differentiable communication. In previous work, Jaques et al. (2018) propose a loss to influence other agents but require explicit and complex models of other agents and their experiments focus on mixed cooperative-competitive scenarios. Eccles et al. (2019) add biases to each agent’s loss function that separately encourage positive listening (i.e., the listener to act differently for different messages) and positive signaling (i.e., the speaker to produce diverse messages in different situations). Their method is simpler but requires task-specific hyperparameter tuning to achieve reasonable performance and underperforms in sensory-rich environments (Lin et al., 2021). Our work is closest to Lin et al. (2021), who leverage autoencoding as their method to learn a message protocol in cooperative 2D MARL games. Agents learn to reconstruct their observations and communicate their autoencoding. It outperforms previous works while being algorithmically and conceptually simpler. Our method builds on this encoding perspective by considering other agents’ messages to ground communication. Whereas agents in Lin et al. (2021) can only learn to encode the observation, our approach leverages the relationship between different agents’ messages to encode global state information. Empirically, our method is also more efficient as it requires no extra learning parameters whereas Lin et al. (2021) learn and discard their decoder network. Note that our setup uses continuous messages instead of discrete (Eccles et al., 2019; Lin et al., 2021), a standard choice in contrastive learning (Chopra et al., 2005; He et al., 2020; Chen et al., 2020a) and embodied multi-agent communication (Sukhbaatar et al., 2016; Jiang and Lu, 2018; Das et al., 2019).
@@ 3 Preliminaries
We base our investigations on decentralized partially observable Markov decision processes (Dec-POMDPs) with N agents to describe a fully cooperative multi-agent task (Oliehoek and Amato, 2016). A Dec-POMDP consists of a tuple $G=\langle S,A,P,R,Z,\Omega,n,\gamma\rangle$. $s\in S$ is the true state of the environment. At each time step, each agent $i\in N$ chooses an action $a\in A^{i}$ to form a joint action $a\in A\equiv A^{1}\times A^{2}...\times A^{N}$. It leads to an environment transition according to the transition function $P(s^{\prime}|s,a^{1},...a^{N}):S\times A\times S\rightarrow[0,1]$. All agents share the same reward function $R(s,a):S\times A\rightarrow\mathbb{R}$. $\gamma\in[0,1)$ is a discount factor. As the environment is partially observable, each agent $i$ receives individual observations $z\in Z$ based on the observation function $\Omega^{i}(s):S\rightarrow Z$.
@@
To account for communication, similar to Lin et al. (2021), at each time step $t$, an agent $i$ takes an action $a_{t}^{i}$ and produces a message $m_{t}^{i}=\Psi^{i}(\Omega^{i}(s_{t}))$ after receiving its observation $\Omega^{i}(s_{t})$ and messages sent at the previous time step $m_{t-1}^{-1}$, where $\Psi^{i}$ is agent $i$’s function to produce a message given its observation and $m_{t-1}^{-1}$ refers to messages sent by agents other than agent $i$. The messages are continuous vectors of dimensionality $D$.
@@ 4 Methodology
We propose a different perspective on the message space used for communication. At each time step $t$ for a given trajectory $\tau$, a message $m_{t}^{i}$ of an agent $i$ can be viewed as an incomplete view of the environment state $s_{t}$ because $m_{t}^{i}$ is a function of $s_{t}$ as formulated in section 3. Naturally, messages of all the N agents are different incomplete perspectives of $s_{t}$. To ground decentralized communication, we hypothesize that we could leverage this relationship between messages from similar states to encourage consistency and similarity of the messages space across agents. Specifically, we propose maximizing the mutual information using contrastive learning which aligns the message space by pushing messages from similar states closer together and messages of different states further apart. Note that agents see a partial view of the state from their observation, so they will inherently communicate different messages to reflect their partial knowledge. However, aligning their message space enables them to communicate the specific parts of the state they observe in a more mutually-intelligible way.
@@
Let $M$ be all the messages in a batch of trajectories and $M_{\tau}$ be the messages in trajectory $\tau$. Let $m^{i}_{t}\in M_{\tau}$ the message of agent $i$ at time $t$. Thus, positives $P$ for a message $m_{t}^{i}$ given a timestep window $w$ are all other messages from the same trajectory $\tau$ sent within that timestep window $P(m_{t}^{i})\equiv\{m_{t^{\prime}}^{j}\in M_{\tau}\setminus\{m_{t}^{i}\}:t^{\prime}\in[t-w,t+w]\}$. Let all other messages $A$ from all trajectories in the batch be $A{(m_{t}^{i})}\equiv M\setminus\{m_{t}^{i}\}$. Formally, the contrastive loss $L_{CACL}$:
$\sum_{m_{t}^{i}\in M}\frac{-1}{\lvert P(m_{t}^{i})\rvert}\sum_{m_{p}\in P(m_{t}^{i})}\log\frac{\exp(m_{t}^{i}\cdot m_{p}/\eta)}{\sum_{m_{a}\in A(m_{t}^{i})}\exp(m_{t}^{i}\cdot m_{a}/\eta)}$ (1)
Where $\eta\in\mathbb{R^{+}}$ is a scalar temperature and $|P(m_{t}^{i})|$ is the cardinality. Practically, each agent has a replay buffer that maintains a batch of trajectory data collected from multiple environment instances. It contains messages received during training to compute the CACL loss. We use a timestep window of size 5 for all the environments based on hyperparameter tuning of different window sizes. Following Khosla et al. (2020), messages are normalized before the loss computation and a low temperature (i.e. $\eta=0.1$) is used as it empirically benefits performance and training stability. The total loss for each agent is a reinforcement learning loss $L_{RL}$ using the reward to learn a policy (but not message head) and a separate contrastive loss $L_{CACL}$ to learn just the message head, formulated as follows:
@@
where $\kappa\in\mathbb{R^{+}}$ is a hyperparameter to scale the CACL loss.
@@ 5.1 Experimental Setup
We evaluate our method on three multi-agent environments with communication channels. Given the limited information each agent observes themselves, agents must meaningfully communicate in order to improve task performance.
Traffic-Junction: Proposed by Sukhbaatar et al. (2016), it consists of A 4-way traffic junction with cars entering and leaving the grid. The goal is to avoid collision when crossing the junction. We use 5 agents with a vision of 1. Although not necessary, given the limited vision in agents, communication could help in solving the task. We evaluate each algorithm with success rate during evaluation episodes. All results are averaged over 12 evaluation episodes and over 6 random seeds. More details of the environments and parameters can be found in appendix A.1.
@@
Find-Goal: Proposed by Lin et al. (2021), agents’ goal is to reach the green goal location as fast as possible in a grid environment with obstacles. We use 3 agents, each observes a partial view of the environment centered at its current position. Unlike in Lin et al. (2021), we use a field of view of $3\times 3$ instead of $5\times 5$ to make the problem harder. Each agent receives an individual reward of 1 for reaching the goal and an additional reward of 5 when all of them reach the goal. Hence, it is beneficial for an agent to communicate the goal location once it observes the goal. As in Lin et al. (2021), we measure performance using episode length. An episode ends quicker if agents can communicate goal locations to each other more efficiently. Hence, a method performs better if it has shorter episode lengths.
@@ 5.3 Task Performance
Figure 2: CACL (red) outperforms all other methods on Traffic-Junction (left), Predator-Prey (left) and Find-Goal (right). Predator-Prey shows evaluation reward, higher is better. Traffic-Junction plots the percent of successful episodes, higher is better. Find-Goal plots the episode length until the goal is reached, lower is better. The performance curves are smoothed by a factor of 0.5 with standard errors are plotted as shaded areas.
@@
We confirm the effectiveness of CACL with an ablation study of the key design decisions: sliding window and SupCon. CACL leverages the temporal nature of RL to treat a sliding window of timesteps as positive views of each other. We plot results for a range of window sizes run on Predator-Prey in Figure 4. No sliding window (size 1) performs poorly demonstrating its necessity and that the choice of sliding window size is an important hyperparameter. Through the use of SupCon (Khosla et al., 2020) we treat all sent and received messages in the sliding window as all positive views of each other, with many positives per batch. Creating a batch with just one positive view per message corresponds to SimCLR (Chen et al., 2020a) and results in much worse performance ($1.36\pm 9.46$). We also run Predator-Prey and search across values of the CACL loss coefficient $\kappa$ in Figure 4. We used the best values (5-step window, $\kappa$=0.5) across all the environments, demonstrating that the choice of CACL hyperparameters is robust. Overall, we show the issues in naively implementing contrastive learning for communication, and the clear, important design decisions behind CACL.
@@ 5.5 Protocol Symmetry
We hypothesize that CACL’s improved performance over the baselines is because it induces a more consistent, mutually-intelligible language that is shared among agents. More specifically, we consider a language’s consistency to be how similarly agents communicate (i.e., sending similar messages) when faced with the same observations. A consistent protocol can reduce the optimization complexity since agents only need to learn one protocol for the whole group and it also makes agents more mutually intelligible.
@@
To evaluate consistency, we measure protocol symmetry (Graesser et al., 2019) so if an agent swaps observations and trajectory with another agent, it should produce a similar message as what the other agent produced. We extend this metric from previous work to the continuous, embodied case and measure the pairwise cosine similarities of messages sent by different agents for the same observation. Let $\binom{N}{k}$ denote the set of all $k$-agent subset of a set of $N$ agents. Given a trajectory $\tau$ and $\{t\in T\}$ as a set of time steps of $\tau$, protocol symmetry ($protocol\_sym$) is written as:
$\frac{1}{|T|}\sum_{i\in T}\frac{1}{|N|}\sum_{i\in N}\frac{1}{|{\binom{N}{k}}|}\sum_{j,k\in{\binom{N}{k}}}\frac{m_{j}\cdot m_{j}}{\|m_{j}\|\|m_{k}\|}$ (3)
Therefore, a more consistent protocol has higher symmetry. We swap agent trajectory and observations and compute this metric over 10 evaluation episodes for 6 random seeds, and show results in Table 1. The self-supervised methods (CACL and AEComm) clearly outperform the others (DIAL and PL) implying that SSL is better for learning consistent representations in decentralized MARL. Furthermore, CACL’s protocol is very highly symmetric, clearly outperforming all others. Each AEComm agent autoencodes their own observation without considering the other agents’ messages, leading to the formation of multiple protocols between agents. In contrast, CACL induces a common protocol by casting the problem in the multi-view perspective and implicitly aligning agents’ messages. The possible correlation between protocol symmetry and overall performance and speed further indicates the benefits of learning a common language in the decentralized setting.

### v3

# Learning Multi-Agent Communication with Contrastive Learning

@@ Abstract
Figure 1: In multi-view learning, augmentations of the original image or “views” are used as positive samples for contrastive learning. In our proposed method, CACL, different agents’ views of the same environment state are considered positive samples and messages are contrastively learned as encodings of that state.
@@ 1 Introduction
Communication is a key capability necessary for effective coordination among agents in partially observable environments. In multi-agent reinforcement learning (MARL) (Sutton & Barto, 2018), agents can use their actions to transmit information (Grupen et al., 2020) but continuous or discrete messages on a communication channel (Foerster et al., 2016), i.e., linguistic communication (Lazaridou & Baroni, 2020), are more flexible and powerful because they can convey more complex concepts. To successfully communicate, a speaker and a listener must share a common language with a shared understanding of the symbols being used (Skyrms, 2010; Dafoe et al., 2020). Emergent communication or learning a common protocol (Wagner et al., 2003; Lazaridou & Baroni, 2020), is a thriving research direction but most works focus on simple, single-turn, sender-receiver games (Lazaridou et al., 2018; Chaabouni et al., 2019). In more visually and structurally complex MARL environments (Samvelyan et al., 2019), existing approaches often rely on centralized learning mechanisms by sharing models (Lowe et al., 2017) or gradients (Sukhbaatar et al., 2016).
@@
Inspired by the literature in representation learning that uses different “views” of a data sample (Bachman et al., 2019), for a given trajectory, we frame an agent’s observation as a “view” of the environment state. Thus, different agents’ messages are encodings of different incomplete “views” of the same underlying state. From this perspective, messages from the same state should, generally, be more similar to each other than to those from distant states or other trajectories, as shown in Figure 1. As with image augmentations, two agent observations may not necessarily overlap, but a contrastive objective can generally lead to effective encodings. We propose Communication Alignment Contrastive Learning (CACL), where each agent separately uses contrastive learning between their own sent and received messages to learn a communication encoding.
@@ 2 Related Work
Learning to coordinate multiple RL agents is a challenging and unsolved task where naively applying single-agent RL algorithms often fails (Foerster et al., 2016). Recent approaches focus on neural network-based agents (Goodfellow et al., 2016) with a message channel to develop a common communication protocol (Lazaridou & Baroni, 2020). To handle issues of non-stationarity, some work focuses on centralized learning approaches that globally share models (Foerster et al., 2016), training procedures (Lowe et al., 2017), or gradients (Sukhbaatar et al., 2016) among agents. This improves coordination and can reduce optimization issues but results are often still sub-optimal in practise (Foerster et al., 2016; Lin et al., 2021) and may violate independence assumptions, effectively modelling the multi-agent scenario as a single agent (Eccles et al., 2019).
This work focuses on independent, decentralized agents and non-differentiable communication. In previous work, Jaques et al. (2018) propose a loss to influence other agents but require explicit and complex models of other agents and their experiments focus on mixed cooperative-competitive scenarios. Eccles et al. (2019) add biases to each agent’s loss function that separately encourage positive listening (i.e., the listener to act differently for different messages) and positive signaling (i.e., the speaker to produce diverse messages in different situations). Their method is simpler but requires task-specific hyperparameter tuning to achieve reasonable performance and underperforms in sensory-rich environments (Lin et al., 2021). Our work is closest to Lin et al. (2021), who leverage autoencoding as their method to learn a message protocol in cooperative 2D MARL games. Agents learn to reconstruct their observations and communicate their autoencoding. It outperforms previous works while being algorithmically and conceptually simpler. Our method builds on this encoding perspective by considering other agents’ messages to ground communication. Whereas agents in Lin et al. (2021) can only learn to encode the observation, our approach leverages the relationship between different agents’ messages to encode global state information. Empirically, our method is also more efficient as it requires no extra learning parameters whereas Lin et al. (2021) learn and discard their decoder network. Note that our setup uses continuous messages instead of discrete (Eccles et al., 2019; Lin et al., 2021), a standard choice in contrastive learning (Chopra et al., 2005; He et al., 2020; Chen et al., 2020a) and embodied multi-agent communication (Sukhbaatar et al., 2016; Jiang & Lu, 2018; Das et al., 2019).
@@ 3 Preliminaries
We base our investigations on decentralized partially observable Markov decision processes (Dec-POMDPs) with N agents to describe a fully cooperative multi-agent task (Oliehoek & Amato, 2016). A Dec-POMDP consists of a tuple $G=\langle S,A,P,R,Z,\Omega,n,\gamma\rangle$. $s\in S$ is the true state of the environment. At each time step, each agent $i\in N$ chooses an action $a^{i}\in A^{i}$ to form a joint action $a\in A\equiv A^{1}\times A^{2}...\times A^{N}$. It leads to an environment transition according to the transition function $P(s^{\prime}|s,a^{1},...a^{N}):S\times A\times S\rightarrow[0,1]$. All agents share the same reward function $R(s,a):S\times A\rightarrow\mathbb{R}$. $\gamma\in[0,1)$ is a discount factor. As the environment is partially observable, each agent $i$ receives individual observations $z\in Z$ based on the observation function $\Omega^{i}(s):S\rightarrow Z$.
@@
To account for communication, similar to Lin et al. (2021), at each time step $t$, an agent $i$ takes an action $a_{t}^{i}$ and produces a message $m_{t}^{i}=\Psi^{i}(\Omega^{i}(s_{t}))$ after receiving its observation $\Omega^{i}(s_{t})$ including messages from the previous time step $m_{t-1}^{-i}$, where $\Psi^{i}$ is agent $i$’s function to produce a message given its observation and $m_{t-1}^{-i}$ refers to messages sent by agents other than agent $i$. The messages are continuous vectors of dimensionality $D$.
@@ 4 Methodology
We propose a different perspective on the message space used for communication. At each time step $t$ for a given trajectory $\tau$, a message $m_{t}^{i}$ of an agent $i$ can be viewed as an incomplete view of the environment state $s_{t}$ because $m_{t}^{i}$ is a function of $s_{t}$ as formulated in section 3. Naturally, messages of all the N agents are different incomplete perspectives of $s_{t}$. To ground decentralized communication, we hypothesize that we could leverage this relationship between messages from similar states to encourage consistency and similarity of the messages space across agents. Specifically, we propose maximizing the mutual information using contrastive learning which aligns the message space by pushing messages from similar states closer together and messages of different states further apart. Note that agents see a partial view of the state from their observation, so they will inherently communicate different messages to reflect their partial knowledge. However, aligning their message spaces enables communicating these partial views of the state in a more mutually-intelligible way.
@@
Let $M$ be all the messages in a batch of trajectories and $M_{\tau}$ be the messages in trajectory $\tau$. Let $m^{i}_{t}\in M_{\tau}$ the message of agent $i$ at time $t$. Thus, positives $H$ for a message $m_{t}^{i}$ given a timestep window $w$ are all other messages from the same trajectory $\tau$ sent within that timestep window $H(m_{t}^{i})\equiv\{m_{t^{\prime}}^{j}\in M_{\tau}\setminus\{m_{t}^{i}\}:t^{\prime}\in[t-w,t+w]\}$. Let all other messages $K$ from all trajectories in the batch be $K{(m_{t}^{i})}\equiv M\setminus\{m_{t}^{i}\}$. Formally, the contrastive loss $L_{CACL}$:
$\sum_{m_{t}^{i}\in M_{\tau}}\frac{-1}{\lvert H(m_{t}^{i})\rvert}\sum_{m_{h}\in H(m_{t}^{i})}\log\frac{\exp(m_{t}^{i}\cdot m_{h}/\eta)}{\sum_{m_{k}\in K(m_{t}^{i})}\exp(m_{t}^{i}\cdot m_{k}/\eta)}$ (1)
Where $\eta\in\mathbb{R^{+}}$ is a scalar temperature and $|H(m_{t}^{i})|$ is the cardinality. Practically, each agent has a replay buffer that maintains a batch of trajectory data collected from multiple environment instances. It contains messages received during training to compute the CACL loss. We use a timestep window of size 5 for all the environments based on hyperparameter tuning of different window sizes. Following Khosla et al. (2020), messages are normalized before the loss computation and a low temperature (i.e. $\eta=0.1$) is used as it empirically benefits performance and training stability. The total loss for each agent is a reinforcement learning loss $L_{RL}$ using the reward to learn a policy (but not message head) and a separate contrastive loss $L_{CACL}$ to learn just the message head, formulated as follows:
@@
where $\kappa\in\mathbb{R^{+}}$ is a hyperparameter to scale the CACL loss. See Appendix A.1 for $L_{RL}$ details.
@@ 5.1 Experimental Setup
We evaluate our method on three multi-agent environments with communication channels. Given the limited information each agent observes, agents must meaningfully communicate in order to improve task performance. All results are averaged over 12 evaluation episodes and over 6 random seeds. More details of the environments and parameters can be found in appendix A.2.
Traffic-Junction: Proposed by Sukhbaatar et al. (2016), it consists of a 4-way traffic junction with cars entering and leaving the grid. The goal is to avoid collision when crossing the junction. We use 5 agents with a vision of 1. Although not necessary, given the limited vision in agents, communication could help in solving the task. We evaluate algorithms using their success rate in avoiding all collisions during evaluation episodes.
@@
Find-Goal: Proposed by Lin et al. (2021), agents’ goal is to reach the green goal location as fast as possible in a grid environment with obstacles. We use 3 agents, each observes a partial view of the environment centered at its current position. Unlike in Lin et al. (2021), we use a field of view of $3\times 3$ instead of $7\times 7$ to make the problem harder. Each agent receives an individual reward of 1 for reaching the goal and an additional reward of 5 when all of them reach the goal. Hence, it is beneficial for an agent to communicate the goal location once it observes the goal. As in Lin et al. (2021), we measure performance using episode length. An episode ends quicker if agents can communicate goal locations to each other more efficiently. Hence, a method performs better if it has shorter episode lengths.
@@ 5.3 Task Performance
Figure 2: CACL (red) outperforms all other methods on Traffic-Junction (left), Predator-Prey (left) and Find-Goal (right). Predator-Prey shows evaluation reward, higher is better. Traffic-Junction plots the percent of successful episodes, higher is better. Find-Goal plots the episode length until the goal is reached, lower is better. The performance curves are smoothed by a factor of 0.5 with standard errors plotted as shaded areas.
@@
We confirm the effectiveness of CACL with an ablation study of the key design decisions: sliding window and SupCon. CACL leverages the temporal nature of RL to treat a sliding window of timesteps as positive views of each other. We plot results for a range of window sizes run on Predator-Prey in Figure 4. No sliding window (size 1) performs poorly, demonstrating its necessity and that the choice of sliding window size is an important hyperparameter. Through the use of SupCon (Khosla et al., 2020) we treat all sent and received messages in the sliding window as all positive views of each other, with many positives per batch. Creating a batch with just one positive view per message corresponds to SimCLR (Chen et al., 2020a) and results in much worse performance ($1.36\pm 9.46$). We also run Predator-Prey and search across values of the CACL loss coefficient $\kappa$ in Figure 4. We used the best values (5-step window, $\kappa$=0.5) across all the environments, demonstrating that the choice of CACL hyperparameters is robust. Overall, we show the issues in naively implementing contrastive learning for communication, and the clear, important design decisions behind CACL.
@@ 5.5 Protocol Symmetry
We hypothesize that CACL’s improved performance over the baselines is because it induces a more consistent, mutually-intelligible communication protocol that is shared among agents. More specifically, we consider consistency to be how similarly agents communicate (i.e., sending similar messages) when faced with the same observations. A consistent protocol can reduce the optimization complexity since agents only need to learn one protocol for the whole group and it also makes agents more mutually intelligible.
@@
To evaluate consistency, we measure protocol symmetry (Graesser et al., 2019) so if an agent swaps observations and trajectory with another agent, it should produce a similar message as what the other agent produced. We extend this metric from previous work to the continuous, embodied case. We feed the same trajectory to all agents and measure the pairwise cosine similarities of the messages that they produce. Given a trajectory $\tau$ and $\{t\in T\}$ as a set of time steps of $\tau$, protocol symmetry ($protocol\_sym$) is written as:
$\frac{1}{|T|}\sum_{t\in T}\frac{1}{|N|}\sum_{i\in N}\frac{1}{|N|-1}\sum_{j\in N\setminus i}\frac{\Psi^{i}(\Omega^{j}(s_{t}))\cdot\Psi^{j}(\Omega^{j}(s_{t}))}{\|\Psi^{i}(\Omega^{j}(s_{t}))\|\|\Psi^{j}(\Omega^{j}(s_{t}))\|}$ (3)
Therefore, a more consistent protocol has higher symmetry. We swap agent trajectory and observations and compute this metric over 10 evaluation episodes for 6 random seeds, and show results in Table 1. The self-supervised methods (CACL and AEComm) clearly outperform the others (DIAL and PL) implying that SSL is better for learning consistent representations in decentralized MARL. Furthermore, CACL’s protocol is very highly symmetric, clearly outperforming all others. Each AEComm agent autoencodes their own observation without considering the other agents’ messages, leading to the formation of multiple protocols between agents. In contrast, CACL induces a common protocol by casting the problem in the multi-view perspective and implicitly aligning agents’ messages. The possible correlation between protocol symmetry and overall performance and speed further indicates the benefits of learning a common protocol in the decentralized setting.